# Continual Learning of Large Language Models: A Comprehensive Survey

### v2

@@ 4.2.2. Different Domains of DAP
The three aforementioned works highlight the importance of DAP for code LLMs. However, it is crucial to note that the problem definition and conventional architectures of existing Code LLMs may present challenges of compatibility for DAP deployment, and need to be addressed in the future.
@@ 4.3. Continual Fine-Tuning (CFT)
Background of Continual Fine-Tuning (CFT). Continual Fine-Tuning (CFT) lies at the bottom layer of the vertical continuity, where models are trained on successive homogeneous tasks drawn from an evolving data distribution. As the service-oriented layer of LLM, it doesn’t require consideration of further adaptation to another downstream tasks, simplifying optimization objectives to a great extent: better adaptation and less forgetting [^3] We direct interested readers to additional survey literature on the topic of general CFT (biesialska2020continual, ; ke2023continual, ).. In the era of LLMs, new computational paradigms in CFT have emerged and attracted significant attention within the research community. These topics include:
  • Continual Instruction Tuning (CIT), where models must generalize to new tasks encoded in instructions, requiring semantic understanding (zhang2023citb, ) (Section 4.3.3).
  • Continual Model Refinement (CMR), where fine-grained, possibly example-level solutions are required, differing from task-level approaches (hartvigsen2023aging, ) (Section 4.3.4).
  • Continual Model Alignment (CMA), which aligns models with evolving human preferences, challenging due to subjective nature and lack of clear task boundaries (lin2024mitigating, ; zhangcppo, ) (Section 4.3.5).
  • Continual Learning for Multimodal Language Models (CMLLMs), where addressing the composite architectural design and preventing catastrophic forgetting are key challenges (he2023continual, ; ni2023continual, ) (Section 4.3.6).
We summarize existing studies on CFT in Table 3, categorizing studies into sub-categories as listed above. The table includes details on incremental learning types (X-IL), LLM architecture, and employed CL techniques and evaluation metrics. After discussing general observations on CFT in Section 4.3.1, we’ll delve into each sub-category in detail.
@@ 4.3.1. General Observations on CFT
  • OBS-1: There has been a noticeable transition in focus from CIL to TIL and DIL. It has been a longstanding common sense in the CL community that CIL, as it requires the model to predict the context label and within-context label at the same time (van2022three, ; wang2024comprehensive, ; kim2022theoretical, ), is the most challenging CL scenario and hence receives most of the attention from the community. However, among all 35 papers presented in Table 3, only 3 papers study CFT of CIL. The transition of the research focus demonstrates the importance of TIL and DIL in the real-world applications of continual LLMs. More detail discussion of this transition will be included in Section 6.2.
@@ 4.3.2. General Continual Fine-Tuning (General CFT)
Researchers have long investigated the phenomenon of forgetting resilience in pre-trained LLMs when fine-tuned for downstream tasks (ke2021achieve, ; tao2022can, ; luo2023investigating, ; zheng2023learn, ; mehta2023empirical, ), despite some discover the opposite (luo2023investigating, ). Although the pre-trained weights initially position the model in a flat-loss basin, aiding adaptation to future tasks without severely impacting previous ones (hao2019visualizing, ; neyshabur2020being, ; mirzadeh2022wide, ; mehta2023empirical, ), zero or near-zero forgetting is only observed at the representation level. This implies that while the model retains its ability to distinguish between task-specific representations, it may still forget specific task details (wu2021pretrained, ; tao2022can, ; luo2023investigating, ; zheng2023learn, ). Therefore, additional measures are necessary when deploying these models in real-world applications (ke2021achieve, ; wei2022circle, ; bai2023enhancing, ; qin2021lfpt5, ; weyssow2023usage, ; chen2024parameterizing, ).
@@
Another line of approaches refines the input/output format and network architectures of pre-trained LLMs to be better suited for CFT. For instance, CTR (ke2021achieve, ) incorporates two CL-plugin modules that consist of a task-specific module (TSM) for acquiring task-specific knowledge and a knowledge-sharing module (KSM) for selectively transferring previously learned similar knowledge. CIRCLE (wei2022circle, ) manually designs diverse prompt templates for various types of buggy code, unifying them as the cloze task and employs difficulty-based replay to enhance continual program repair. LFPT5 (qin2021lfpt5, ) addresses lifelong few-shot language learning by consolidating sequence labeling, text classification, and text generation into a text-to-text generation task. It undergoes prompt tuning on generated pseudo-examples from previous domains when adapting to new tasks. In (zhang2022continual, ), the authors propose a method for adaptively adding compositional adapters during continual sequence generation tasks. Before training on new domains, a decision stage determines which trained module can be reused. During training, this module also regenerates examples of the past for replay. C3 (chen2024parameterizing, ) merges PEFT and in-context learning (ICL) in a teacher-student framework. The teacher model undergoes in-context tuning focused solely on the current domain, while the student model, together with tunable prompts, minimizes the KL-divergence between the output distribution and the ground truth and teacher model simultaneously.
@@ 4.3.3. Continual Instruction Tuning (CIT)
While LLMs are typically pre-trained on extensive and diverse corpora, they may struggle with specific tasks such as instruction following despite their general knowledge. Numerous studies have shown that Instruction Tuning (IT) can notably improve LLMs’ ability to follow textual instructions (zhang2024instruction, ; wei2021finetuned, ; jiang2024instructiontuned, ; sanh2022multitask, ; ouyang2022rlhf, ), leveraging the pre-existing knowledge within LLMs to bridge the gap between general and task-specific performance (wei2022finetuned, ). Recent works like WizardLM (xu2023wizardlm, ) and CodecLM (wang2024codeclm, ) further tailor synthetic data to steer LLMs’ behavior through IT. Additionally, IT enhances the interaction between humans and LLMs, providing a more natural interface and aligning LLM outputs more closely with human expectations and preferences (luo2023empirical, ).
@@ 4.3.4. Continual Model Refinement (CMR)
Like humans, LLMs are prone to errors, such as inaccurate translations or outdated information (de2021editing, ). Directly fine-tuning the model to correct these mistakes may disrupt its performance on previously learned tasks. To overcome these challenges, Model Refinement (also known as Model Editing) is proposed, which aims to rectify the model’s errors while preserving its performance on other inputs, with only moderate computing resources (sinitsin2020editable, ; de2021editing, ; fast_edit, ; hase2021language, ; huang2023transformer, ; mitchell2022memory, ; hartvigsen2023aging, ). The concept of model editing was initially explored in (sinitsin2020editable, ), which introduced a “reliability-locality-efficiency” principle and proposed a gradient descent editor to address it efficiently. Subsequent research, such as (de2021editing, ) and (fast_edit, ), extended this principle to edit factual knowledge in BERT-based language models and larger models like GPT-J-6B (gpt-j, ) and T5-XXL (raffel2020exploring, ), respectively, using gradient decomposition. These approaches typically update a subset of model parameters to alter the labels of specific inputs. Additionally, memory-based models, as discussed in (mitchell2022memory, ) and (hartvigsen2023aging, ), incorporate editing through retrieval mechanisms.
The concept of Continual Model Refinement (CMR) extends model refinement horizontally, presenting updated sample pairs ${({\bm{x}}_{e},y_{e},\widehat{y}_{e})}^{e=1}_{N}$ sequentially as a stream. (lin2022continual, ) initially introduces this idea, evaluating various CL methods with a dynamic sampling algorithm. Many CMR methods employ a retrieval mechanism. For instance, (hartvigsen2023aging, ) uses hidden activations of the language model as a “key” to activate updated parameters only when input $x_{0}$ resembles updated sample pairs; (yu2023melo, ) improves this approach’s efficiency by integrating LoRA (hu2021lora, ); (das2024larimar, ) augments the LLM with an external episodic memory, modeling CMR as an ongoing memory refresh. Meanwhile, some methods focus solely on updating a subset of model parameters. For example, (hu2024wilke, ) addresses the issue of “toxicity buildup and flash” in single-editing methods like ROME (meng2022locating, ), adapting it to the CL context with a knowledge-aware layer selection algorithm. WISE (wang2024wise, ) addresses the “impossible triangle” of reliability, locality, and generalization in existing lifelong model refinement methods. It introduces a side memory system that enables knowledge sharding and merging, successfully achieving all three objectives simultaneously.
@@ 4.3.5. Continual Model Alignment (CMA)
Model Alignment (MA) ensures AI systems’ actions and outputs align with human values, ethics, and preferences (ouyang2022rlhf, ; rafailov2024dpo, ). MA can be broadly categorized into two types: Reinforcement Learning-based (RL-based) and Supervised Learning-based (SL-based). RL-based approaches (wu2022survey, ; ouyang2022rlhf, ; schulman2017proximal, ) are trained to make decisions reinforced by human feedback, using a reward system to guide them towards desirable outcomes. Conversely, the SL-based approaches (hendrycks2023aligning, ; rafailov2024dpo, ; ji2024ai, ) directly train models on datasets of human preferences, aligning their output with demonstrated human values. Both approaches leverage a combination of algorithmic learning techniques and human feedback to progressively refine the model behavior. When LLMs undergo the phase of MA, vertical forgetting of previous knowledge usually occurs. In (lin2024mitigating, ), the authors refer to this phenomenon of catastrophic forgetting induced caused by MA as the “Alignment Tax”. Notably, even a single stage of MA can diminish the model’s performance capabilities, as it restricts the model’s responses to a narrower subset of the training distribution.
Continual Model Alignment (CMA) aims to continuously refine LLMs to align with evolving human values, ethics, and data. The static nature of LLM training on historical data sets can lead to discrepancies between the models’ outputs and current factual accuracies, societal norms, and standards, making CMA a crucial process for maintaining their adaptability and alignment with contemporary contexts (taori2023alpaca, ). Likewise, there are two types of CMA frameworks: RL-based and SL-based. In the realm of RL-based CMA, two significant contributions have been noted. (lin2024mitigating, ) identifies the conflicts between the existing CL techniques and RLHF, and proposes Adaptive Model Averaging (AMA), adaptively finding appropriate ratios for the combination of model layers to gain maximal rewards with minimal tax; Continual Proximal Policy Optimization (CPPO) (zhangcppo, ) proposes a weighting strategy for different examples deciding its usage of policy enhancement or knowledge retention, mitigating the alignment tax over time. For SL-based CMA, Continual Optimal Policy Fitting (COPF) (zhang2023copf, ) presents a solution adapted from the Direct Policy Optimization (DPO) (rafailov2024direct, ), solving its potential risks of sub-optimal policy fitting and over-optimization in the context of CMA.
@@ 4.3.6. Continual Multimodal Large Language Models (CMLLMs)
Multi-modal LLMs integrate data of multiple modalities, like texts, images and videos to enhance real-world information comprehension (peng2023kosmos2, ; li2023otter, ). Typically, MLLMs consist of modality-specific sub-modules such as pre-trained vision encoders, large language models, and projectors for cross-model alignment. This alignment is essential for MLLMs to fuse the diverse data types and promote their comprehension.
Continually training multi-modal models like CLIP (radford2021learning, ) has been long studied (zheng2023preventing, ; ni2023continual, ; wu2024building, ; khan2021personalizing, ; liu2023class, ; garg2024tic, ; li2024coleclip, ; jha2024clap4clip, ; yu2024select, ; yu2024boosting, ), while the problem of continually training MLLMs still remains underexplored. Several existing studies have investigated the causes of catastrophic forgetting when continually training MLLMs. (zheng2024antiforgetting, ) performs singular value decomposition on input embeddings, revealing a significant disparity among different input embeddings. This discrepancy causes the model to learn irrelevant information for previously trained tasks, resulting in catastrophic forgetting and negative forward transfer. (zhai2023investigating, ) observes that minority collapse may lead to catastrophic forgetting, when the imbalance ratio between majority and minority classes approaches infinity during fine-tuning. It further identifies hallucination as a contributing factor to performance degradation in MLLMs.
@@ 5. Evaluation Protocols and Datasets
In this section, we introduce the evaluation protocols and datasets for continul LLMs. In Section 5.1, we discuss common continual learning evaluation metrics adapted for this context, along with metrics designed specifically for continual LLMs. Then, in Section 5.2, we outline the datasets available for each discussed topic.
5.1. Evaluation Protocols
5.1.1. Interpreting Basic Continual Learning Evaluation Metrics in the Continual LLMs
In some literature, $\widetilde{\operatorname{OP}}$ is referred to as “example accuracy” (chen2024parameterizing, ), “whole accuracy” (song2023conpet, ), or “edit success rate” in CMR (hartvigsen2023aging, ). The concepts of Forgetting and Backward Transfer underpin various evaluation metrics, such as knowledge retention (jin2022lifelong, ), performance on unchanged knowledge (jang2022temporalwiki, ), average increased perplexity (AP+(qin2022elle, ), and test and edit retention rate in CMR (hartvigsen2023aging, ). We extend the notation of forward transfer in the vertical direction to represent the performance improvement on downstream tasks resulting from domain-adaptive pre-training (see Table 2). Forward Transfer is alternatively referred to as temporal generalization (jin2022lifelong, ) or knowledge transfer (lazaridou2021mind, ) in some literature.
5.1.2. Evaluation Metrics of Continual LLMs
(14) $\Delta R_{t}^{X}$ $\triangleq\tfrac{1}{M}\sum_{m=1}^{M}(R_{t,i}^{X}-b_{i}^{X}).$
5.2. Datasets

### v3

@@ 4.2.2. Different Domains of DAP
The three aforementioned studies highlight the importance of DAP for code LLMs. However, it is crucial to note that the problem definition and conventional architectures of existing Code LLMs may present challenges of compatibility for DAP deployment, and need to be addressed in the future.
@@ 4.3. Continual Fine-Tuning (CFT)
Background of Continual Fine-Tuning (CFT). Continual Fine-Tuning (CFT) lies at the bottom layer of the vertical continuity, where models are trained on successive homogeneous tasks drawn from an evolving data distribution. As the service-oriented layer of LLM, it does not require consideration of further adaptation to another downstream tasks, simplifying optimization objectives to a great extent: better adaptation and less forgetting [^2] We direct interested readers to additional survey literature on the topic of general CFT (biesialska2020continual, ; ke2023continual, ).. In the era of LLMs, new computational paradigms in CFT have emerged and attracted significant attention within the research community. These topics include (i) Continual Instruction Tuning (CIT) (zhang2023citb, ), (ii) Continual Model Refinement (CMR) (hartvigsen2023aging, ), (iii) Continual Model Alignment (CMA) (lin2024mitigating, ; zhangcppo, ), and (iv) Continual Learning for Multimodal Language Models (CMLLMs) (he2023continual, ; ni2023continual, ). We summarize existing studies on CFT in Table 3, categorizing studies into sub-categories as listed above. The table includes details on incremental learning types (X-IL), LLM architecture, and employed CL techniques and evaluation metrics. After discussing general observations on CFT in Section 4.3.1, we will delve into each sub-category in detail.
@@ 4.3.1. General Observations on CFT
  • OBS-1: There has been a noticeable transition in focus from CIL to TIL and DIL. It has been a longstanding common sense in the CL community that CIL, as it requires the model to predict the context label and within-context label at the same time (van2022three, ; wang2024comprehensive, ; kim2022theoretical, ), is the most challenging CL scenario and hence receives most of the attention from the community. However, among all 35 papers presented in Table 3, only 3 papers study CFT of CIL. The transition of the research focus demonstrates the importance of TIL and DIL in the real-world applications of continual LLMs. More detailed discussion of this transition is included in Section 6.2.
@@ 4.3.2. General Continual Fine-Tuning (General CFT)
Researchers have long investigated the phenomenon of forgetting resilience in pre-trained LLMs when fine-tuned for downstream tasks (ke2021achieve, ; tao2022can, ; luo2023investigating, ; zheng2023learn, ; mehta2023empirical, ), despite some discover the opposite (luo2023investigating, ). Although the pre-trained weights initially position the model in a flat-loss basin, aiding adaptation to future tasks without severely impacting previous ones (mehta2023empirical, ), zero or near-zero forgetting is only observed at the representation level. This implies that while the model retains its ability to distinguish between task-specific representations, it may still forget specific task details (wu2021pretrained, ; tao2022can, ; luo2023investigating, ; zheng2023learn, ). Therefore, additional measures are necessary when deploying these models in real-world applications (ke2021achieve, ; wei2022circle, ; bai2023enhancing, ; qin2021lfpt5, ; weyssow2023usage, ; chen2024parameterizing, ).
@@
Another line of approaches refines the input/output format and network architectures of pre-trained LLMs to be better suited for CFT. For instance, CTR (ke2021achieve, ) incorporates two CL-plugin modules, i.e., a task-specific module (TSM) for acquiring task-specific knowledge and a knowledge-sharing module (KSM) for selectively transferring previously learned similar knowledge. CIRCLE (wei2022circle, ) manually designs diverse prompt templates for various types of buggy code, unifying them as the cloze task and employs difficulty-based replay to enhance continual program repair. LFPT5 (qin2021lfpt5, ) addresses lifelong few-shot language learning by consolidating sequence labeling, text classification, and text generation into a text-to-text generation task. It undergoes prompt tuning on generated pseudo-examples from previous domains when adapting to new tasks. In (zhang2022continual, ), the authors propose a method for adaptively adding compositional adapters during continual sequence generation tasks. Before training on new domains, a decision stage determines which trained module can be reused. During training, this module also regenerates examples of the past for replay. C3 (chen2024parameterizing, ) merges PEFT and in-context learning (ICL) in a teacher-student framework. The teacher model undergoes in-context tuning focused solely on the current domain, while the student model, together with tunable prompts, minimizes the KL-divergence between the output distribution and the ground truth and teacher model simultaneously.
@@ 4.3.4. Continual Model Refinement (CMR)
The concept of model editing was initially explored in (sinitsin2020editable, ), which introduced a “reliability-locality-efficiency” principle and proposed a gradient descent editor to address it efficiently. Subsequent research, such as (de2021editing, ) and (fast_edit, ), extended this principle to edit factual knowledge in BERT-based language models and larger models like GPT-J-6B (gpt-j, ) and T5-XXL (raffel2020exploring, ), respectively, using gradient decomposition. These approaches typically update a subset of model parameters to alter the labels of specific inputs. Additionally, memory-based models, as discussed in (mitchell2022memory, ) and (hartvigsen2023aging, ), incorporate editing through retrieval mechanisms.
Continual Model Refinement (CMR) extends model refinement horizontally, presenting updated sample pairs ${({\bm{x}}_{e},y_{e},\widehat{y}_{e})}^{e=1}_{N}$ sequentially as a stream. (lin2022continual, ) initially introduces this idea, evaluating various CL methods with a dynamic sampling algorithm. Many CMR methods employ a retrieval mechanism. For instance, (hartvigsen2023aging, ) uses hidden activations of the language model as a “key” to activate updated parameters only when input $x_{0}$ resembles updated sample pairs; (yu2023melo, ) improves this approach’s efficiency by integrating LoRA (hu2021lora, ); (das2024larimar, ) augments the LLM with an external episodic memory, modeling CMR as an ongoing memory refresh. Meanwhile, some methods focus solely on updating a subset of model parameters. For example, (hu2024wilke, ) addresses the issue of “toxicity buildup and flash” in single-editing methods like ROME (meng2022locating, ), adapting it to the CL context with a knowledge-aware layer selection algorithm. WISE (wang2024wise, ) addresses the “impossible triangle” of reliability, locality, and generalization in existing lifelong model refinement methods. It introduces a side memory system that enables knowledge sharding and merging, successfully achieving all three objectives simultaneously.
@@ 4.3.5. Continual Model Alignment (CMA)
When LLMs undergo the phase of MA, vertical forgetting of previous knowledge usually occurs. In (lin2024mitigating, ), the authors refer to this phenomenon of catastrophic forgetting induced caused by MA as the “Alignment Tax.” Notably, even a single stage of MA can diminish the model’s performance capabilities, as it restricts the model’s responses to a narrower subset of the training distribution.
Continual Model Alignment (CMA) aims to continuously refine LLMs to align with evolving human values, ethics, and data. The static nature of LLM training on historical data sets can lead to discrepancies between the models’ outputs and current factual accuracies, societal norms, and standards, making CMA a crucial process for maintaining their adaptability and alignment with contemporary contexts. Likewise, there are two types of CMA frameworks: RL-based and SL-based. In the realm of RL-based CMA, two significant contributions have been noted. (lin2024mitigating, ) identifies the conflicts between the existing CL techniques and RLHF, and proposes Adaptive Model Averaging (AMA), adaptively finding appropriate ratios for the combination of model layers to gain maximal rewards with minimal tax; Continual Proximal Policy Optimization (CPPO) (zhangcppo, ) proposes a weighting strategy for different examples deciding its usage of policy enhancement or knowledge retention, mitigating the alignment tax over time. For SL-based CMA, Continual Optimal Policy Fitting (COPF) (zhang2023copf, ) presents a solution adapted from the Direct Policy Optimization (DPO) (rafailov2024direct, ), solving its potential risks of sub-optimal policy fitting and over-optimization in the context of CMA.
@@ 4.3.6. Continual Multimodal Large Language Models (CMLLMs)
Continually training multi-modal models like CLIP (radford2021learning, ) has been long studied (zheng2023preventing, ; ni2023continual, ), while the problem of continually training MLLMs still remains underexplored. Several existing studies have investigated the causes of catastrophic forgetting when continually training MLLMs. (zheng2024antiforgetting, ) performs singular value decomposition on input embeddings, revealing a significant disparity among different input embeddings. This discrepancy causes the model to learn irrelevant information for previously trained tasks, resulting in catastrophic forgetting and negative forward transfer. (zhai2023investigating, ) observes that minority collapse may lead to catastrophic forgetting, when the imbalance ratio between majority and minority classes approaches infinity during fine-tuning. It further identifies hallucination as a contributing factor to performance degradation in MLLMs.
@@ 5. Evaluation Protocols and Datasets
Continual LLMs’ Evaluation Protocols. LAnguage Model Analysis (LAMA) is an evaluation framework designed to probe the world knowledge embedded in language models (petroni2019language, ). LAMA converts each world fact into a cloze statement, which is then input into the language models to predict the correct answer. It has been extensively utilized in work on CPT under the temporal shifts (jang2022temporalwiki, ; jang2022towards, ). FUAR (Forgotten / (Updated + Acquired) Ratio) is proposed for CPT to address the OP’s drawback of not able to accurately reflect the model’s behavior. A FUAR value of 1 represents an equal trade-off between the knowledge forgetting and knowledge learning, while a FUAR less than 1 suggests high learning efficacy. In TRACE (wang2023trace, ), the authors propose a set of “X-Delta” metrics for continual instruction tuning, quantifying the forward transfer on specific abilities of LLMs, which is a straightforward extension of FWT. Specifically, the authors construct three sets of evaluation tasks to benchmark the ability of LLMs, including general ability, instruction following, and safety. For more detailed introduction to these evaluation protocols, please refer to Appendix B.2.
Datasets. In this section, we provide a comprehensive review of the datasets available for benchmarking continual LLMs, as illustrated in Table 4. We provide information about these datasets’ types, what distributional shifts and semantic domains they include, and their sources and applications. We intentionally exclude datasets used for domain-adaptive pre-training LLMs in vertical domains such as legal, medical, and financial, unless they are specifically designed for continual domain-adaptive pre-training. Furthermore, we omit datasets used in general continual fine-tuning, as they have already been extensively studied in existing works (biesialska2020continual, ; ke2023continual, ). For details, please refer to Appendix B.3.